8.0 L of gas at 100°C has a temperature of 88.0 kPa. What will the new temperature of this guys be in Kelvin if the volume is changed to 5.4 L and the pressure changes to 65.3 kPa?


1
Expert's answer
2021-03-29T05:15:58-0400

Using state equation,

P1V1/T1 = P2V2/T2

88.0×8.0/373 = 65.3×5.4/T2

T2 = 186.8 K

= 187 K

So, new temperature = [187 K]